# wd1: Weighted Policy Optimization for Reasoning in Diffusion Language Models

### v1

@@ Abstract
Improving the reasoning capabilities of diffusion-based large language models (dLLMs) through reinforcement learning (RL) remains an open problem. The intractability of dLLMs likelihood function necessitates approximating the current, old, and reference policy likelihoods at each policy optimization step. This reliance introduces additional computational overhead and lead to potentially large bias – particularly when approximation errors occur in the denominator of policy ratios used for importance sampling. To mitigate these issues, we introduce wd1, a novel policy optimization approach that reformulates the objective as a weighted likelihood, requiring only a single approximation for the current parametrized policy likelihood. Experiments on widely used reasoning benchmarks demonstrate that wd1, without supervised fine-tuning (SFT) or any supervised data, outperforms existing RL methods for dLLMs, achieving up to 16% higher accuracy. wd1 delivers additional computational gains, including reduced training time and fewer function evaluations (NFEs) per gradient step. These findings, combined with the simplicity of method’s implementation and R1-Zero-like training (no SFT), position wd1 as a more effective and efficient method for applying RL to dLLMs reasoning.
Code: https://github.com/xiaohangt/wd1
@@ 1 Introduction
Figure 1: Performance on popular reasoning and planning benchmarks with the same base model LLaDA. For all models, we evaluate with maximum length $256$ and $512$, and report the best results. Our method wd1, outperforms both our reproduction of d1 and accuracies of LLaDA (Zhao et al., 2025).
@@
A key challenge in applying reinforcement learning (RL) to dLLMs is the intractability of their likelihood functions (Zhao et al., 2025; Yang et al., 2025), which necessitates approximating the sequence-level likelihoods for policy optimization. Methods such as GRPO (Shao et al., 2024) require estimating the likelihoods of the current, old, and reference policies at every update step, leading to significant computational overhead – especially when large sample sizes are needed for low-error approximations. Moreover, small errors in likelihood estimation can still induce large bias in the objective due to the use of importance sampling, where the importance weights are policy ratios. The inaccuracies in the denominator of policy ratios can severely distort the value of ratio. These issues can be further exacerbated as the completion length and the number of diffusion steps increase.
To address these challenges, we propose wd1, a policy optimization approach with weighted log-likelihood objective for dLLMs. Crucially, this objective dispenses with explicit policy ratios and relies on a single likelihood estimate, thereby avoiding the potentially large bias from policy ratio and reducing the computational overhead. The principal contributions of this work are:
  • We first propose to approximate the closed-form solution of single-iterate reverse-KL-constrained policy optimization in each iteration by optimizing a weighted log-likelihood objective. Crucially, this objective removes the need for policy ratios and depends only on the current policy’s likelihood. We further prove that the resulting iterative updates guarantee monotonic policy improvement.
  • Based on this insight, we introduce wd1, a simple weighted policy optimization method for diffusion language models, where the weights are dependent on the group-relative advantage $\hat{A}_{i}$:
    Our method aims to increase the probability of completions with higher $\hat{A}_{i}$ values and decrease it for lower ones, reinforcing beneficial outcomes and discouraging detrimental ones.
  • In our experiments, we demonstrate that wd1, even without supervised fine-tuning (SFT), outperforms the existing diffusion-based RL method d1 across a suite of reasoning benchmarks, achieving up to a 16% improvement in test accuracy (see Figure 1). wd1 delivers additional speed-ups in RL training, underscoring its effectiveness and establishing it as a substantially more efficient approach for reasoning dLLMs.
@@ 2 Preliminaries
We denote the generation distribution of diffusion-based Large Language Models (dLLMs) by $\pi_{\theta}$. Denote prompt $q\in\mathcal{D}$, and completions $o$. Notably, the reward function denoted by $R(q,o)$ in this report is not limited to verifier. We use superscript $k$ to indicate the $k$-th token in the sequence of completion: $o^{k}$.
@@ 2.1 Diffusion Large Language Models
The most promising discrete diffusion for language modeling is masked diffusion (MD), which gradually corrupt text sequence with mask token (Sahoo et al., 2024; Ou et al., 2025; Shi et al., 2025; Lou et al., 2024). Let $t$ denote the diffusion timestep, and $x_{t}$ the intermediate noised sequence at step $t$. The fully denoised sequence (i.e., the model output) is represented by $x_{0}$, and the forward process is a continuous-time Markov chain with transition denoted by $p_{t|0}(x_{t}\mid x_{0})$.
This work aims to apply RL to LLaDA (Ou et al., 2025; Zhu et al., 2025a), which directly models the clean data distribution conditional on masked sequence $\pi_{\theta}(x_{0}^{k}\mid x_{t})$. By assuming absorbing transition (Austin et al., 2023; Campbell et al., 2022), the Evidence Lower Bound (ELBO) is reduced to a simple objective that has been used for pretraining (Nie et al., 2025b; Ou et al., 2025):
$\max_{\theta}\mathbb{E}_{t\sim\mathcal{U}[0,1],\ x_{0}\sim p_{\text{data}},\ x_{t}\sim p_{t|0}(x_{t}|x_{0})}\left[\frac{1}{t}\sum_{k=1}^{|o_{i}|}\mathbf{1}[x_{t}^{k}=\texttt{mask}]\log\pi_{\theta}(x_{0}^{k}\mid x_{t})\right],$ (1)
where $L$ is the length of the sequence and $x_{0}^{k}$ is the $k$-th token of $x_{0}$. Specifically, intermediate steps $t$ are sampled from uniform distribution, and masked sequence is sampled following the predefined forward process $p_{t|0}(x_{t}\mid x_{0})$. The ELBO can be estimated by the sum of the conditional log-probabilities $\log\pi_{\theta}(x_{0}^{k}\mid x_{t})$ of the tokens that are masked at $t$. This ELBO has also been used to approximate the marginal likelihood $\log\pi_{\theta}(x_{0})$ for fine-tuning (Nie et al., 2025a; Yang et al., 2025; Zhu et al., 2025a).
@@ 2.2 Existing Policy Optimization Methods
The base method of current prevailing RL fine-tuning algorithms are Trust Region Policy Optimization (TRPO) (Schulman et al., 2015) or called Proximal Policy Optimization (PPO) with fixed KL penalty (Schulman et al., 2017), in which forward KL (FKL) divergence is applied to restrict the policy update:
@@
where $A^{\pi_{\text{old}}}$ is the advantage function, $q$ and $o$ are denoted as the prompt and (clean) response, respectively. Theorem 1 (see Appendix A) demonstrates the monotonic policy improvement of FKL regularized policy optimization.
@@
PPO then extends the soft constraint (KL penalty) to clipping policy ratio $\pi_{\theta}(\cdot|q)/\pi_{\text{old}}(\cdot|q)$ for restricting the policy update, further employed in fine-tuning (Ouyang et al., 2022) with additional reverse-KL regularization w.r.t. the reference policy $\pi_{\text{ref}}$. Group Relative Policy Optimization (GRPO) (Shao et al., 2024) simplifies PPO by sampling a group of completions $o_{1:G}$ and approximating their advantage with their normalized rewards. This advantage is corrected by removing the standard deviation (Liu et al., 2025) $\hat{A}_{i}=R(q,o_{i})-\texttt{mean}(R(q,o_{1:G}))$, which is called group-relative advantage. Denote $r_{i}^{k}$ to be policy ratio $\pi_{\theta}(o^{k}_{i})/\pi_{\text{old}}(o^{k}_{i})$, and for conditional distribution $a,b$, $D^{\mathcal{D}}_{\text{KL}}(a\|b)=\mathbb{E}_{\begin{subarray}{c}q\sim\mathcal{D}\end{subarray}}[D_{\text{KL}}(a(\cdot|q)\|b(\cdot|q))]$, the GRPO objective is
$\mathcal{L}_{\text{GRPO}}(\theta)=\mathbb{E}_{\begin{subarray}{c}q\sim\mathcal{D},\\ \,o_{1:G}\sim\pi_{\theta}(\cdot\mid q)\end{subarray}}\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\min\Big{(}r_{i}^{k}\hat{A}_{i},\,\text{clip}(r_{i}^{k},1-\varepsilon,1+\varepsilon)\hat{A}_{i}\Big{)}\Bigg{]}-\beta D^{\mathcal{D}}_{\text{KL}}\Big{[}\pi_{\theta}\|\pi_{\text{ref}}\Big{]}.$ (3)
2.3 Group Relative Policy Optimization for Masked dLLMs
@@
Adapting GRPO to diffusion-based large language models (dLLMs) presents notable challenges, primarily due to the method’s dependence on sequence-level log-likelihoods. Unlike autoregressive models, dLLMs generate outputs via a non-autoregressive, iterative denoising process, making the computation of $\log\pi_{\theta}(o|q)$ intractable and necessitating approximation.
@@
Existing works LLaDA (Nie et al., 2025a) and MMaDA (Yang et al., 2025) employ ELBO as the per-token log-likelihood approximation $\phi^{\pi}(o^{k})=\phi^{\pi}(x_{0}^{k})=\exp(\mathbb{E}_{t\in\mathcal{U}[0,1]}[\mathbf{1}[x_{t}^{k}=\mathtt{mask}]\log\pi(x_{0}^{k}|x_{t},q)])$. However, an accurate estimation requires a large sample size, resulting in inefficiency for online RL. An alternative method introduced in d1 (Zhao et al., 2025) is $\phi^{\pi}(o^{k})=\phi^{\pi}(x_{0}^{k})=\pi(x_{0}^{k}|q^{\prime})$, where $q^{\prime}$ is randomly masked prompt. To obtain the sequence-level log-likelihood, all the current approaches are based on the mean-field approximation $\log\pi(o|q)\approx\sum_{k}\log\pi(x_{0}^{k}|q)$.
Diffusion-based GRPO (Zhao et al., 2025; Yang et al., 2025) can then be written as:
@@
$\mathbb{E}_{q\sim\mathcal{D},\,o_{1:G}\sim\pi_{\text{old}}(\cdot\mid q)}\Bigg{[}\tfrac{1}{G}\sum_{i=1}^{G}\tfrac{1}{|o_{i}|}\sum_{k=1}^{|o_{i}|}\min\bigg{(}\tfrac{\phi^{\pi_{\theta}}(o_{i}^{k})}{\phi^{\pi_{\text{old}}}(o_{i}^{k})},1-\epsilon,1+\epsilon\bigg{)}\hat{A}_{i}-\beta D_{\text{KL}}\big{[}\phi^{\pi_{\theta}}(\cdot)\,\|\,\phi^{\pi_{\text{ref}}}(\cdot)\big{]}\Bigg{]}.$ (4)
However, regardless of the specific approximation method employed, the approximated likelihood appears in the denominator of the policy ratio, where even minor estimation errors can introduce disproportionately large bias when $\phi^{\pi_{\theta}}(o_{i}^{k})>\phi^{\pi_{\text{old}}}(o_{i}^{k})$, and thus lead to instability during training. Moreover, the approximation function $\phi$ is typically applied to three separate policies $\pi_{\theta}$, $\pi_{\text{old}}$, and $\pi_{\text{ref}}$ – resulting in additional computational cost compared to RL for AR models. In this report, we seek a policy optimization method that minimizes reliance on approximated likelihoods, thereby improving both fine-tuning stability and efficiency for diffusion-based large language models.
@@
3.1 Objective Derivation
As shown in previous works (Peng et al., 2019; Rafailov et al., 2023), policy optimization with reverse-KL regularization has a closed form solution. Approximating the target solution does not rely on importance sampling, and thus eliminates the dependence on policy ratio, i.e., $\pi_{\theta}(\cdot|q)/\pi_{\text{old}}(\cdot|q)$, and extensive likelihood approximation. Therefore, we propose to use reverse-KL regularization with respect to the old policy, along with standard reference policy regularization:
@@ 3.1 Objective Derivation
Notably, reverse-KL regularized policy optimization objective in Equation 5 also enjoys monotonic improvement guarantees, as established in Theorem 3.
We note that the two reverse KL divergences in Equation 5 can be merged as $(\lambda+\beta)D_{\text{KL}}\big{(}\pi_{\theta}(\cdot|q)\,\|\,\pi^{\text{ref}}_{{\text{old}}}(\cdot|q)\big{)}$. Here $\pi_{\text{old}}^{\text{ref}}(\cdot|q)\propto\pi_{\text{old}}(\cdot|q)^{\lambda/(\lambda+\beta)}\cdot\pi_{{\text{ref}}}(\cdot|q)^{\beta/(\lambda+\beta)}$, is a geometric mixture between the reference and the old policy. We further recognize the solution to Equation 5 has the following form (Peng et al., 2019):
$\pi^{*}(\cdot|q)\propto\pi_{{\text{old}}}^{\text{ref}}(\cdot|q)\cdot\exp\big{(}\psi A^{\pi_{\text{old}}}(q,\cdot)\big{)},$ (6)
where we let $\psi:={1}/{(\lambda+\beta)}$. Then, we can instead optimize the KL divergence $D_{\text{KL}}(\pi^{*}(\cdot|q)\,\|\,\pi_{\theta}(\cdot|q))$, resulting in the negative log-likelihood (NLL) loss $\mathcal{L}_{\text{NLL}}(\theta):=-\mathbb{E}_{q\sim\mathcal{D},o\sim\pi^{*}}[\log\pi_{\theta}(o|q)]$. We apply importance sampling with the proposal distribution $\pi_{\text{old}}^{\text{ref}}(o|q)$ and importance weights proportional to $\exp\big{(}\psi A^{\pi_{\text{old}}}(q,o)\big{)}$ to approximate the NLL loss (see Theorem 2):
@@
$\mathcal{L}_{\text{NLL}}(\theta)\approx-\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{\mathrm{G}}\sim\pi_{\text{old}}^{\text{ref}}(\cdot|q)}\Big{[}\sum_{i=1}^{\mathrm{G}}\frac{\exp\big{(}\psi A^{\pi_{\text{old}}}(q,o_{i})\big{)}}{\sum_{j=1}^{\mathrm{G}}\exp\big{(}\psi A^{\pi_{\text{old}}}(q,o_{j})\big{)}}\cdot\log\pi_{\theta}(o_{i}|q)\Big{]}$ (7)
Therefore, we obtain an alternative objective in Equation 7 that does not involve the policy ratio $\pi_{\theta}(\cdot|q)/\pi_{\text{old}}(\cdot|q)$ or $\pi_{\theta}(\cdot|q)/\pi_{\text{ref}}(\cdot|q)$. In contrast to the approaches of Zhao et al. (2025) and Zhu et al. (2025a), our algorithm only needs to approximate $\pi_{\theta}(\cdot\mid q)$.
3.2 Method

### v2

@@ Abstract
Improving the reasoning capabilities of diffusion-based large language models (dLLMs) through reinforcement learning (RL) remains an open problem. The intractability of dLLMs likelihood function necessitates approximating the current, old, and reference policy likelihoods at each policy optimization step. This reliance introduces additional computational overhead, and can lead to large variance and estimation error in RL objective – particularly in computing the policy ratio for importance sampling. To mitigate these issues, we introduce wd1, a novel ratio-free policy optimization approach that reformulates the RL objective as a weighted log-likelihood, requiring only a single approximation for the current parametrized policy likelihood. We formally show that our proposed method can be interpreted as energy-guided discrete diffusion training combined with negative sample unlearning, thereby confirming its theoretical soundness. In experiments on LLaDA-8B model, wd1 outperforms diffusion-based GRPO (d1) while requiring lower computational cost, achieving up to a $59\%$ improvement in accuracy. Furthermore, we extend wd1 to denoising-stepwise weighted policy optimization (wd1++), achieving state-of-the-art math performance of $44.2\%$ on MATH500 and $84.5\%$ on GSM8K with only 20 RL training steps.
@@ 1 Introduction
A key challenge in applying reinforcement learning (RL) to dLLMs is the intractability of their likelihood functions (Zhao et al., 2025; Yang et al., 2025), which necessitates approximation for policy optimization. Applying approximated log-likelihood for diffusion-based GRPO (Shao et al., 2024; Zhao et al., 2025) can exponentially amplify the approximation error and lead to large variance when computing the policy ratio for importance sampling. Moreover, GRPO requires the estimated likelihoods of the current, old, and reference policies at every training step, leading to significant computational overhead. These issues can be further exacerbated as the completion length and the number of diffusion steps increase.
To address these challenges, we propose wd1, a policy optimization approach with weighted log-likelihood objective for dLLMs. Crucially, this objective dispenses with explicit policy ratios and relies on a single likelihood approximation, thereby avoiding the potentially large bias and variance in policy ratio, and reducing the computational overhead. Our principal contributions are:
  • We propose a novel reinforcement learning method for dLLMs, termed wd1, which formulates the objective as a weighted log-likelihood of outcome sequence, derived from reverse-KL regularized policy optimization. The weight, defined as $(-w^{+}+w^{-})$ and dependent on the advantage $A$, balances two terms: $w^{+}\propto\exp(A)$ increases the probability of higher-advantage completions, while $w^{-}\propto\exp(-A)$ decreases the probability of lower-advantage ones. Together, this mechanism amplifies beneficial outcomes meanwhile actively reducing detrimental ones.
  • We prove that our proposed RL method for dLLMs can be interpreted as jointly training an energy-guided discrete diffusion model—guided by the advantage function—and unlearning low-advantage data, thereby steering generation toward higher-advantage completions.
  • We conduct experiment with LLaDA-8B-Instruct model (Nie et al., 2025a). Compared to the baseline method d1 (Zhao et al., 2025), our method wd1 achieves 76.4% on Sudoku (Arel, 2025) (+58.8% over d1) and 51.2% on Countdown (Pan et al., 2025) (+16% over d1), without requiring supervised fine-tuning (SFT), and with significantly less computational burden in RL training.
  • We further extend our method to leverage intermediate completions generated in the decoding process, which we call wd1++. The extended method surpasses several concurrent RL for dLLMs methods, achieving state-of-the-art performance 44.2% on MATH500 and 84.5% on GSM8K with only 20 training steps, and $10\times$ fewer rollouts compared to the baseline methods.
@@ 2 Preliminaries
We denote the generation policy of diffusion-based Large Language Models (dLLMs) by $\pi_{\theta}$. Denote prompt $q\in\mathcal{D}$, and completions $o\in\mathcal{O}$. Notably, the reward function denoted by $R(q,o)$ in this work is not limited to verifiers. We use superscript $k$ to indicate the $k$-th token of completion: $o^{k}$ or $x_{0}^{k}$.
@@ 2.1 Diffusion Large Language Models
The most promising discrete diffusion for language modeling is masked diffusion, which gradually corrupts the text sequence with a mask token (Sahoo et al., 2024; Ou et al., 2025; Shi et al., 2025; Lou et al., 2024). Let $t\in[0,1]$ denote the diffusion timestep, and $x_{t}$ as the masked sequence at step $t$. The fully denoised sequence (i.e., the completion $o$) is represented by $x_{0}$, and the forward process ($p_{t|0}(x_{t}\mid x_{0})$) is formulated as a continuous-time Markov chain. The transition kernel $\mathbf{Q}_{t}$ is absorbing (Austin et al., 2023; Campbell et al., 2022), meaning that at time $t$, $\mathbf{Q}_{t}=\sigma(t)\mathbf{Q}^{\text{absorb}}$, where $\sigma$ is a decreasing scalar noise schedule and $\mathbf{Q}^{\text{absorb}}$ is a constant matrix (See Definition 2).
This work aims to apply reinforcement learning to fine-tune masked discrete diffusion models such as LLaDA (Ou et al., 2025; Zhu et al., 2025a), which models the clean data distribution conditional on masked sequence as $\pi_{\theta}(x_{0}^{k}\mid x_{t})$. The negative Evidence Lower Bound (ELBO) is reduced to a simple objective called Denoising Cross Entropy (DCE) (Ou et al., 2025): $\forall x_{0}\sim p_{\text{data}}$,
$\mathcal{L}_{\text{DCE}}(x_{0})=-\mathbb{E}_{t\sim\mathcal{U}[0,1],\ x_{t}\sim p_{t|0}(x_{t}|x_{0})}\left[\frac{1}{t}\sum_{k=1}^{K}\mathbf{1}(x_{t}^{k}=[\texttt{mask}])\log\pi_{\theta}(x_{0}^{k}\mid x_{t})\right],$ (1)
where $K$ is the length of the sequence and $x_{0}^{k}$ is the $k$-th token of $x_{0}$. Specifically, intermediate steps $t$ are sampled from uniform distribution, and masked sequence is sampled following the predefined forward process $p_{t|0}(x_{t}\mid x_{0})$. DCE can be used to approximate the marginal likelihood $\log\pi_{\theta}(x_{0})$ for supervised fine-tuning and reinforcement learning (Nie et al., 2025a; Yang et al., 2025).
@@ 2.2 Existing Policy Optimization Methods
The base method of current prevailing RL fine-tuning algorithms is Trust Region Policy Optimization (TRPO) (Schulman et al., 2015), in which forward KL divergence is applied to restrict the update:
@@
where $A^{\pi_{\text{old}}}$ is the advantage function, $q$ and $o$ are denoted as the prompt and (clean) response, respectively. Proposition˜1 (Appendix˜A) demonstrates the monotonic policy improvement of TRPO.
@@
PPO then extends the soft constraint (KL penalty) to clipping policy ratio $\pi_{\theta}(\cdot|q)/\pi_{\text{old}}(\cdot|q)$ and employing pessimism for policy update, further employed in fine-tuning (Ouyang et al., 2022) with additional reverse-KL regularization w.r.t. the reference policy $\pi_{\text{ref}}$. Group Relative Policy Optimization (GRPO) (Shao et al., 2024) simplifies PPO by sampling a group of completions $\{o_{i}\}_{i=1}^{G}$ and approximating their advantage with their normalized rewards. This advantage is corrected by subtracting the mean reward across the group (Liu et al., 2025): $\hat{A}_{i}=R(q,o_{i})-\texttt{mean}\big(R(q,o_{1:G})\big),$ which we refer to as the group-relative advantage.
2.3 Policy Optimization for dLLMs
@@
Adapting GRPO to diffusion-based large language models (dLLMs) presents notable challenges, since dLLMs generate outputs via a non-autoregressive, iterative denoising process, making the computation of $\log\pi_{\theta}(o|q)$ intractable and necessitating approximation for policy optimization.
@@
Existing works by Nie et al. (2025a); Yang et al. (2025) employ ELBO for per-token log-likelihood approximation following DCE: $\phi^{\pi}(x_{0}^{k})=\mathbb{E}_{t\in\mathcal{U}[0,1]}[w\cdot\mathbf{1}[x_{t}^{k}=\mathtt{mask}]\log\pi(x_{0}^{k}|x_{t},q)]$, where $w=1/t$ in DCE and $w=1$ in UniGRPO (Yang et al., 2025). However, an accurate estimation requires a large sample size of $t$, resulting in inefficiency for online RL. A biased but efficient method is introduced in d1 (Zhao et al., 2025), requiring only sample at $t=1$: $\phi^{\pi}(x_{0}^{k})=\log\pi(x_{0}^{k}|x_{1},q^{\prime})$, where prompt $q^{\prime}$ is randomly masked, $x_{1}$ is fully masked response.
In diffusion-based GRPO (Zhao et al., 2025; Yang et al., 2025), policy ratio is then computed using the approximated log-likelihoods: $r^{k}_{i}(\theta)=\pi_{\theta}(o^{k}_{i})/\pi_{\text{old}}(o^{k}_{i})\approx\exp\big(\phi^{\pi_{\theta}}(o_{i}^{k})-\phi^{\pi_{\text{old}}}(o_{i}^{k})\big)$ for importance sampling in estimating the objective of GRPO:
@@
$\mathbb{E}_{\begin{subarray}{c}q\sim\mathcal{D},\\ \,o_{1:G}\sim\pi_{\text{old}}(\cdot\mid q)\end{subarray}}\bigg[\frac{1}{GK}\sum_{i=1}^{G}\sum_{k=1}^{K}\min\big(r^{k}_{i}(\theta)\hat{A}_{i},\text{clip}(r^{k}_{i}(\theta),1\pm\epsilon)\hat{A}_{i}\big)-\beta D_{\text{KL}}\big(\pi_{\theta}(\cdot)\,\|\,\pi_{\text{ref}}(\cdot)\big)\bigg].$ (3)
Figure 1: Example policy ratio value $r_{i}^{k}$ computed using ELBO and approximated likelihood in d1 on GSM8K after a policy update. Ratio’s unclipped interval is $1-\epsilon,1+\epsilon]$, where $\epsilon=0.5$. ELBO-based likelihood approximation yields high-variance ratio estimates; d1 induces a biased ratio that can deviate substantially from ELBO. Both methods suffer from efficiently and accurately compute ratios.
However, existing approaches are hampered by their reliance on extensive likelihood approximation to compute the policy ratio. In current diffusion-based GRPO methods, the ratio is computed as $r_{i}^{k}\approx\exp\big(\phi^{\pi_{\theta}}(o_{i}^{k})-\phi^{\pi_{\text{old}}}(o_{i}^{k})\big)$ so approximation errors in likelihood can be exponentially amplified. As formally shown in Appendix A.1, the resulting error in the estimated RL objective becomes more severe when less accurate log-likelihood approximations are used, such as in d1, or ELBO used in DCE and UniGRPO when the Monte Carlo sample size $t$ is small.
[…]
@@
In this section, we introduce wd1, a novel RL algorithm that eliminates the need for approximating the likelihood (policy) ratios for importance sampling, aiming to reduce the computational burden, and the variance and approximation error in computing the RL objective. We further extend our method to wd1++ by applying denoising-stepwise policy optimization.
3.1 Reinforcement Learning as Weighted Log-Likelihood Maximization
Prevailing RL methods are based on constrained policy optimization (Belousov and Peters, 2017), penalizing the deviation of current policy $\pi_{\theta}(\cdot|q)$ from the old policy $\pi_{\text{old}}(\cdot|q)$. TRPO objective (Equation˜2) applies a forward-KL penalty. We instead adopt reverse-KL penalty augmented with the reference policy regularization $D_{\text{KL}}(\pi_{\theta}(\cdot|q)\,\|\,\pi_{{\text{ref}}}(\cdot|q))$:
@@ 3.1 Reinforcement Learning as Weighted Log-Likelihood Maximization
Note that the monotonic improvement guarantee still holds when using reverse-KL penalty, as we show in Theorem 2. From the method of Lagrange multipliers, the solution to Equation˜4 has the following form (Peng et al., 2019; Rafailov et al., 2023):
$\pi^{*}(\cdot|q)\propto\pi_{\text{old}}(\cdot|q)^{\lambda/(\lambda+\beta)}\cdot\pi_{{\text{ref}}}(\cdot|q)^{\beta/(\lambda+\beta)}\cdot\exp\bigg(\frac{A^{\pi_{\text{old}}}(q,\cdot)}{\lambda+\beta}\bigg).$ (5)
As the analytic form of the optimal policy $\pi^{*}$ is known, we can train our policy by directly minimizing $D_{\text{KL}}(\pi^{*}(\cdot|q)\,\|\,\pi_{\theta}(\cdot|q))$. This minimization can be expressed as the following weighted log-likelihood (WLL) loss, where the weights $\propto\exp\big(\psi A^{\pi_{\text{old}}}\big)$, $\psi=\tfrac{1}{\lambda+\eta}$ and the samples are obtained from the geometric mixture policy $\pi_{\text{old}}^{\text{ref}}(\cdot|q)\propto\pi_{\text{old}}(\cdot|q)^{\lambda/(\lambda+\beta)}\cdot\pi_{{\text{ref}}}(\cdot|q)^{\beta/(\lambda+\beta)}$ (See Proposition˜2): $\forall q\sim\mathcal{D}$,
$\mathcal{L}_{\text{WLL}}(\theta)$ $=\mathbb{E}_{o\sim\pi^{\text{ref}}_{\text{old}}(\cdot|q)}\Big[-\exp\big(\psi A^{\pi_{\text{old}}}(q,o)\big)\cdot\log\pi_{\theta}(o|q)\Big]$ (6)
@@
$\approx\mathbb{E}_{\{o_{i}\}_{i=1}^{G}\sim\pi^{\text{ref}}_{\text{old}}(\cdot|q)}\left[\frac{1}{G}\sum_{i=1}^{G}-\frac{\exp\big(\psi\hat{A}_{i}\big)}{\sum_{j=1}^{G}\exp\big(\psi\hat{A}_{j}\big)}\log\pi_{\theta}(o_{i}|q)\right].$ (7)
As shown in Equation˜7, we approximate the advantage function using the group-relative advantage $\hat{A}$ and normalize the weights, thereby limiting their magnitude and reducing variance in loss computation. Notably, dividing by the normalization constant does not affect the solution, since it is independent of the completions. The resulting objective does not involve ratio $\pi_{\theta}(\cdot|q)/\pi_{\text{old}}(\cdot|q)$ for importance sampling or $\pi_{\theta}(\cdot|q)/\pi_{\text{ref}}(\cdot|q)$ for regularization, successfully avoiding the potential amplification of log-likelihood approximation error and large variance in diffusion GRPO.
Although the objective $\mathcal{L}_{\text{WLL}}(\theta)$ in Equation˜7 avoids the likelihood ratio estimation, it has two limitations. First, the algorithm is not fully utilizing all the completions. Due to the exponential form of the weighting, completions with relatively low advantage – referred to as negative samples – may receive vanishingly small weights, and do not contribute to learning. Second, due to the likelihood-maximization property of WLL, the likelihoods of all sampled completions are increased, even for negative samples. This issue is exacerbated in scenarios where all completions attain identical but low rewards (e.g. 0), thus all weights become equal and the likelihoods of these suboptimal samples are nonetheless reinforced.
3.2 wd1: Fully Utilizing Completions
We propose wd1, an improved weighted log-likelihood objective that explicitly reinforces positive samples and penalizes negative samples:
$\mathcal{L}_{\textit{wd1}}(\theta)=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\text{old}}^{\text{ref}}(\cdot|q)}\Big[\frac{1}{G}\sum_{i=1}^{G}\big(-w^{+}(q,o_{i})+w^{-}(q,o_{i})\big)\cdot\log\pi_{\theta}(o_{i}|q)\Big],$ (8)
where the weights are based on group-relative (GRPO) advantage and are further normalized to avoid overly imbalanced weight $\hat{A}_{i}=R(q,o_{i})-\texttt{mean}(R(q,o_{1:G}))$:
wd1 objective balances positive and negative samples through a complementary penalty term, $w^{-}(q,o_{i})\,\log\pi_{\theta}(o_{i}|q)$, which minimizes the likelihood of low-advantage completions. This penalty induces negative gradients, thereby accelerating divergence from undesirable completions. Moreover, in the extreme case where all completions exhibit identical advantages, the optimization naturally halts since $w^{+}=w^{-}$, thereby addressing the concern on increasing likelihood of negative samples proposed in Sec 3.1. We demonstrate the effectiveness of this combination via ablations in C.2.
Our method wd1, a simple ratio-free policy optimization based on weighted log-likelihood objective for dLLMs, is formally presented in Algorithm˜1. We first obtain $G$ completions $\{o\}_{i=1}^{G}$ sampled from geometric mixture $\pi_{\text{old}}^{\text{ref}}(\cdot|q)\propto\pi_{\text{old}}(\cdot|q)^{\lambda/(\lambda+\beta)}\cdot\pi_{{\text{ref}}}(\cdot|q)^{\beta/(\lambda+\beta)}$ (line 5). Since the base model LLaDA parametrizes the clean token prediction $\pi^{\text{ref}}_{\text{old}}(x_{0}^{k}|x_{t},q)$ for denoising, we approximate $\log\pi^{\text{ref}}_{\text{old}}(x_{0}^{k}|x_{t},q)\approx\lambda\log\pi_{\text{old}}(x_{0}^{k}|x_{t},q)+\beta\log\pi_{\text{ref}}(x_{0}^{k}|x_{t},q)$ as the logits of the denoising distribution at each step $t$. We then use the samples to compute weights in Equation˜9 (line 6). In weights computing, we leverage completions from all groups to estimate the normalization constant, in order to restrict the the gradient norm and stabilize the training. Finally in line 8, we approximate the log-likelihood of completions, and compute objectives for policy update. Likelihood approximation in d1 (Zhao et al., 2025) is directly applicable to wd1: $\log\pi_{\theta}(x_{0}|q)\approx\sum_{k}\log\pi_{\theta}(x_{0}^{k}|x_{1},q^{\prime})$, where $q^{\prime}$ is randomly masked from prompt $q$ at every gradient step.
3.3 wd1++: Stepwise Weighted Policy Optimization
The decoding process in dLLMs relies on confidence-based remasking (Wang et al., 2025b). At each denoising step $l\in\{1,\cdots,L\}$ in decoding, clean data is predicted conditional on the masked sequence $x_{l}$ and then tokens with low-confidence are re-masked for further denoising, which construct a refinement process. Since current diffusion RL methods only use the final predicted clean completion for training, there are bunch of clean completions in the intermediate denoising steps remaining unused.
To leverage intermediate clean completions, we extend our weighted log-likelihood objective to a step-wise formulation based on DCE, which we term wd1++. In wd1 (as well as in GRPO), a group of completions $\{o_{i}\}_{i=1}^{G}$ is sampled for policy optimization. In wd1++, we expand this group to $\{O_{i}\}_{i=1}^{G}$, where $O_{i}=\{x_{0|l}\mid x_{0|l}\sim\pi_{\text{old}}^{\text{ref}}(\cdot\mid x_{t},q),\;x_{0|L}=o_{i},\;l=1,\dots,L\},$ which contains all generated completions during the decoding process, including intermediate ones. The expanded group of completions is then used to estimate both the advantage function and the corresponding weights. The resulting loss objective is defined as:
$\mathcal{L}_{\textit{wd1}++}(\theta)=\mathbb{E}_{\begin{subarray}{c}q\sim\mathcal{D},\\ \{O_{i}\}_{i=1}^{G}\sim\pi_{\text{old}}^{\text{ref}}(\cdot|q)\\ l\in\text{Unif}\{1,\cdots,L\}\end{subarray}}\big[\tfrac{L}{Gl}\sum_{i=1}^{G}\sum_{x_{0|l}\in O_{i}}\big(-w^{+}(q,x_{0|l})+w^{-}(q,x_{0|l})\big)\cdot\log\pi_{\theta}(x_{0|l}|x_{l},q)\big].$ (10)